# SE4Lip: Speech-Lip Encoder for Talking Head Synthesis to Solve Phoneme-Viseme Alignment Ambiguity

### v1

@@ 4.2 Speech Feature Extractor
NeRF [29] Wav2vec 2.0 [3] 31.6359 (-1.17%) 0.0422 (-1.17%) 3.7204 (+25.6%) 3.3782 (-55.9%) 10.200 (+42.8%)
Whisper [31] 31.2311 (-2.44%) 0.0436 (+2.11%) 3.6727 (+24.0%) 3.3416 (-56.4%) 10.202 (+42.8%)
SE4Lip(Ours) 32.2301 (+0.68%) 0.0399 (-6.56%) 2.8725 (-3.01%) 8.7098 (+13.7%) 6.1255 (-14.2%)
@@
3DGS [23] Wav2vec 2.0 [3] 30.8003 (-0.42%) 0.0421 (+2.43%) 3.2513 (+9.72%) 3.9522 (-21.9%) 9.6083 (+4.29%)
Whisper [15] 30.5449 (-0.28%) 0.0410 (-0.24%) 3.1397 (+5.93%) 3.2848 (-35.1%) 10.344 (+12.3%)
SE4Lip(Ours) 31.0176 (+0.29%) 0.0401 (-2.43%) 2.6836 (-9.44%) 8.3487 (+64.8%) 6.6705 (-27.6%)
@@
The extracted STFT spectrogram is processed through a Gated Recurrent Unit (GRU) network. The core of GRU consists of the update gate $z_{t}$ and the reset gate $r_{t}$. For the STFT spectrogram input sequence$\{x_{1},...,x_{T}\}$, the definitions of $z_{t}$ and $r_{t}$ are given in Eq. 1 and Eq. 2.
$z_{t}=\sigma(W_{z}\cdot[h_{t-1},x_{t}])$ (1)
$r_{t}=\sigma(W_{r}\cdot[h_{t-1},x_{t}])$ (2)
Here, $\sigma$ denotes the Sigmoid function, $h_{t-1}$ is the hidden state at the previous time step, and $W_{z}$ and $W_{r}$ are the corresponding weight matrices. The computation process of GRU at time step $t$ is represented by Eq. 3 and Eq. 4.
$\tilde{h}_{t}=\tanh(W_{h}\cdot[r_{t}\odot h_{t-1},x_{t}])$ (3)
$h_{t}=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$ (4)
4.3 Lip Feature Extractor
We aim to provide accurate lip features for aligning speech and lip shapes. Inspired by works [30, 7], we designed a CNN-based model to extract lip features. The specific architecture of this model is shown in Tab. 2. The model captures fine-grained spatial features of the lip images layer by layer through multiple convolutional layers. After processing the lip shape image through the CNN, an embedding is obtained with the same dimension as the speech features, representing the dynamic visual features of the lip shapes.
4.4 Loss Function
We use cosine similarity to measure the similarity between speech embedding and lip shape embedding as Eq. 5.
$cos(a,v)=\frac{a\cdot v}{||a||||v||}$ (5)
Here, $a$ and $v$ represent the speech and lip shape embeddings. For each speech-lip feature pair, we use cross-entropy loss as the contrastive loss function. The loss function encourages correctly paired speech and lip features to be closer to each other in the feature space, while incorrectly paired speech and lip features are pushed farther apart. The specific loss function is shown in Eq. 6.
$\mathcal{L}=-y\cdot\log(cos(a,v))-(1-y)\cdot\log(1-cos(a,v))$ (6)
Here, $y$ is the pairing label, indicating whether the speech and lip shape pair match. In the joint embedding space, the distance between matching speech-lip pairs decreases while the distance between mismatched pairs increases.
5 Experiments
@@ 5.1 Experimental Settings
Dataset. We use the same well-edited video sequences from [29, 23], which contains 6072 frames in total. The video has a frame rate of 25 fps, a resolution of 512x512, and a subject is centered in the video.
Comparison Method. We selected the latest representative works, SyncTalk [29] and TalkingGaussian [23], as rendering models. Specifically, SyncTalk refers to the NeRF-based framework, while TalkingGaussian refers to the 3DGS-based framework. We compare the video quality driven by DeepSpeech [1], HuBERT [15], Wav2Vec2.0 [3], Whisper [31] and SE4Lip.
Implementation Details. SE4Lip uses STFT with hyperparameters is set to n-fft=512, win-length=512, and hop-length=128. The number of GRU layers is set to 8. During training, the learning rate is 5e-5, the window is 5, and the batch size is 16. For SyncTalk, we use SmoothL1 Loss [10] as the loss function. The training steps are set to 60,000, with the fine-tuning steps set to 88,000. For TalkingGaussian, we use L1 Loss as the loss function, and the training steps are set to 20,000. Both training and video rendering are performed on a single NVIDIA RTX 4090 GPU.
@@
SE4Lip(Ours) 2.9304 8.1294 6.9714 2.9362 8.0932 7.0046
Table 4: The quantitative results of lip sync accuracy. We used two different speech samples to synthesize the videos. We highlight the best and second best results.
5.2 Quantitative Evaluation
Evaluation Results. We present the quantitative results in Tab. 3. Our method exhibits significant advantages in lip sync accuracy. Our method’s LMD surpasses the four acoustics features. In the NeRF model, our method improves LSE-C and LSE-D by 13.7% and 14.2%, respectively, compared to the best-performing baseline (HuBERT [15]). This advantage is further amplified in the 3DGS model. Notably, the LSE-C and LSE-D of our method are very close to the ground truth video, indicating the fine-grained capture of speech dynamic information by the STFT-GRU combination and the effectiveness of the cross-modal alignment framework. Our method also achieves the best performance in the visual quality metrics. Additionally, we used out-of-distribution speech to drive video synthesis in the NeRF model. We present the experimental results in Tab. 4. We used LMD, LSE-C, and LSE-D as metrics. The experimental results show that our method outperforms the four acoustics features in terms of lip sync accuracy.
Method DeepSpeech [1] HuBERT [15] Wav2Vec 2.0 [3] Whisper [31] SE4Lip(Ours)
@@ 5.2 Quantitative Evaluation
Table 5: The results of user study. The rating is on a scale of 1-5. A higher scores indicate better performance. We highlight the best and second best results.
@@ 5.3 Qualitative Evaluation
Evaluation Results. To intuitively assess the quality of the synthesized video, we present a comparison between our method and other speech features in Fig. 6. As shown, SE4Lip presents more accurate lip shapes and higher visual quality. In the right half of Fig. 6, we provide a detailed comparison between DeepSpeech [1] and our method. When the movement amplitude of the lip is large, DeepSpeech generates a blurry lip shape (as shown in the first row) and also synthesizes inaccurate lip shape (as shown in the second row). When the lips are fully closed, DeepSpeech reveals the teeth and fails to fully close the lips (as shown in the third row). Our method can closely approach the ground truth video. This is due to the cross-modal alignment framework, which effectively addresses the phoneme-viseme alignment ambiguity issue. And fine-grained speech feature modeling preserves richer feature details. Furthermore, we observe that inaccurate lip features also affect other parts of the image. For example, in the second row of the right half of the figure, the neck area of the subject shows a noticeable shadow when using DeepSpeech.
@@
User Study. We conducted a user study to assess the quality of the synthesized videos effectively. We sampled 20 video clips from the quantitative evaluation and invited 12 volunteers to participate in the study. We used the mean opinion score (MOS) as the metric. Volunteers were asked to rate the synthesized videos on three aspects: 1) Lip-sync Accuracy, 2) Image Quality, and 3) Video Realness. The average scores for each method are presented in Tab. 5. As shown, our method significantly outperforms the comparison methods in terms of Lip-sync Accuracy, indicating the effectiveness of the cross-modal alignment framework.
@@ 5.4 Ablation Study
To evaluate the effectiveness of the SE4Lip framework, we conducted ablation experiments from two dimensions: speech spectrograms and modeling. The experimental results are presented in Tab. 6. As shown, the combination of STFT spectrogram and GRU significantly outperforms other variants in both video fidelity and lip sync accuracy, highlighting the necessity of the collaborative design of modules in SE4Lip.
[…]
@@
5.5 Disscussion
Our experimental results have demonstrated that SE4Lip effectively improves the quality of talking head synthesis, particularly in terms of lips. Although we have implemented SE4Lip using relatively simple models, its effectiveness highlights that the phoneme-viseme alignment ambiguity we discovered is a significant barrier to synthesizing high-quality lips. Additionally, phoneme-viseme alignment ambiguity in other languages also deserves attention.
@@ 6 Conclusion
To address the issue of phoneme-viseme alignment ambiguity in talking head synthesis tasks, we propose the SE4Lip. SE4Lip aligns the speech with the lip shape through a cross-modal alignment framework rather than aligning linguistic representations as in acoustics features. Additionally, SE4Lip processes the speech using a combination of an STFT spectrogram and a GRU-based model. This approach effectively captures fine-grained features in the time and frequency domain, providing strong support for subsequent rendering models. Experimental results show that SE4Lip achieves state-of-the-art performance on both NeRF and 3DGS rendering models. Notably, in terms of lip sync accuracy, SE4Lip improves LSE-C and LSE-D by 13.7% and 14.2%, compared to the best baseline, and produces results close to the ground truth videos. Ablation experiments further present the effectiveness of the STFT-GRU combination.

### v2

@@ 5 Experiments
Wav2vec 2.0 [6] 31.6359 (-1.17%) 0.0422 (-1.17%) 3.7204 (+25.6%) 3.3782 (-55.9%) 10.200 (+42.8%)
NeRF Whisper [7] 31.2311 (-2.44%) 0.0436 (+2.11%) 3.6727 (+24.0%) 3.3416 (-56.4%) 10.202 (+42.8%)
AVHubert [41] 31.8361 (-0.54%) 0.0421 (-1.41%) 2.9273 (-1.16%) 8.0136 (+4.62%) 6.8022 (-4.73%)
Wav2Lip [8] 32.1391 (+0.40%) 0.0406 (-4.92%) 2.9771 (+0.52%) 8.4461 (+10.3%) 6.4028 (-10.3%)
SyncTalk [17] 32.5362(1.33%) 0.0470(-10.1%) 2.8842(-2.61%) 8.6977(11.9%) 6.5323(-8.51%)
PASE(Ours) 32.2301 (+0.68%) 0.0399 (-6.56%) 2.8725 (-3.01%) 8.7098 (+13.7%) 6.1255 (-14.2%)
@@
Wav2vec 2.0 [6] 30.8003 (-0.42%) 0.0421 (+2.43%) 3.2513 (+9.72%) 3.9522 (-21.9%) 9.6083 (+4.29%)
3DGS Whisper [5] 30.5449 (-0.28%) 0.0410 (-0.24%) 3.1397 (+5.93%) 3.2848 (-35.1%) 10.344 (+12.3%)
AVHubert [41] 30.8315 (-0.33%) 0.0408 (-0.73%) 2.8779 (-2.89%) 5.3618 (+5.84%) 8.7027 (-5.53%)
Wav2Lip [8] 30.6718 (-0.85%) 0.0415 (+0.97%) 2.7758 (-6.33%) 6.9973 (+38.1%) 7.7244 (-16.2%)
TalkingGaussian [20] 30.8790(-0.17%) 0.0411(+0%) 3.0011(+1.27%) 4.9713(-1.86%) 9.0187(-2.1%)
PASE(Ours) 31.0176 (+0.29%) 0.0401 (-2.43%) 2.6836 (-9.44%) 8.3487 (+64.8%) 6.6705 (-27.6%)
@@ 5.1 Experimental Settings
Dataset. We use the LRS2 [42] dataset to train PASE. We follow the current works [17, 20, 15, 16] and use the same well-edited video dataset for the face synthesis. This dataset has a frame rate of 25 fps, a resolution of 512x512, and the subject is centered in the video.
Comparison Method. We selected the latest representative works, SyncTalk [17] and TalkingGaussian [20], as rendering models. Specifically, SyncTalk refers to the NeRF-based framework, while TalkingGaussian refers to the 3DGS-based framework. We further replace the speech encoders in both models with DeepSpeech [4], HuBERT [5], Wav2Vec2.0 [6], Whisper [7], AVHubert [41], Wav2Lip [8], and PASE for comparison.
Preprocessing. We train PASE using the LRS2 [42] dataset. Before training, the dataset undergoes a series of preprocessing steps. To obtain accurate phoneme annotations, we employ the Montreal Forced Aligner [43] to align all audio samples in the dataset with corresponding phonemes and timestamps. Based on the alignment results, each video is segmented into phoneme-level audio clips and their associated facial frames. From each audio segment, STFT spectrograms are extracted to serve as the input audio representation. For lip images, we first apply Face Alignment [44] to detect facial landmarks. We then retain only the coordinates corresponding to the lip region (landmarks 48 to 68) to define the lip boundaries. Finally, the lip region is cropped from the full-face image based on the detected landmarks.
Implementation Details. PASE uses STFT with hyperparameters is set to n-fft=512, win-length=512, and hop-length=128. The number of GRU layers is set to 8. During training, the learning rate is 5e-5, the window is 5, and the batch size is 16. For SyncTalk, we use SmoothL1 Loss [45] as the loss function. The training steps are set to 60,000, with the fine-tuning steps set to 88,000. For TalkingGaussian, we use L1 Loss as the loss function, and the training steps are set to 20,000. Both training and video rendering are performed on a single NVIDIA RTX 4090 GPU.
5.2 Quantitative Evaluation
@@
Synctalk [17] 2.9521 8.3814 7.2186 2.9729 7.7299 7.3170
PASE(Ours) 2.9304 8.1294 6.9714 2.9362 8.0932 7.0046
Table 4: The quantitative results of lip accuracy. We highlight the best and second best results.
Evaluation Results. We present the quantitative results in Table 3. Our method exhibits significant advantages in lip sync accuracy. Our method's LMD surpasses the four acoustic features. In the NeRF model, our method improves LSE-C and LSE-D by 13.7% and 14.2%, respectively, compared to the HuBERT [5]. This advantage is further amplified in the 3DGS model. Notably, the LSE-C and LSE-D of our method are close to the ground truth video, indicating the fine-grained capture of speech dynamic information by the STFT-GRU combination and the effectiveness of the cross-modal alignment framework. Furthermore, our approach outperforms state-of-the-art face synthesis methods, including AVHubert [41], Wav2Lip [8], SyncTalk [17], and TalkingGaussian [20]. It is worth noting that SyncTalk [17] employs specially designed speech features, while TalkingGaussian [20] relies on DeepSpeech [4]. We also used out-of-distribution speech to drive video synthesis in the NeRF model. We present the experimental results in Table 4. We used LMD, LSE-C, and LSE-D as metrics. The experimental results show that our method outperforms the four acoustic features in terms of lip sync accuracy.
Method DeepSpeech [4] HuBERT [5] Wav2Vec 2.0 [6] Whisper [7] PASE(Ours)
@@
Table 5: The results of the user study. The rating is on a scale of 1-5. A higher scores indicate better performance. We highlight the best and second best results.
@@ 5.3 Qualitative Evaluation
Evaluation Results. To intuitively assess the quality of the synthesized video, we present a comparison between our method and other speech features in Figure 6. As shown, PASE presents more accurate lip shapes and higher visual quality. In the right half of Figure 6, we provide a detailed comparison between DeepSpeech [4] and our method. When the movement amplitude of the lip is large, DeepSpeech generates a blurry lip shape (as shown in the first row) and also synthesizes an inaccurate lip shape (as shown in the second row). When the lips are fully closed, DeepSpeech reveals the teeth and fails to fully close the lips (as shown in the third row). Our method can closely approach the ground truth video. This is due to the cross-modal alignment framework, which effectively addresses the phoneme-viseme alignment ambiguity issue. And fine-grained speech feature modeling preserves richer feature details. Furthermore, we observe that inaccurate lip features also affect other parts of the image. For example, in the second row of the right half of the figure, the neck area of the subject shows a noticeable shadow when using DeepSpeech.
@@
User Study. We conducted a user study to assess the quality of the synthesized videos effectively. We sampled 20 video clips from the quantitative evaluation and invited 12 volunteers to participate in the study. We used the mean opinion score (MOS) as the metric. Volunteers were asked to rate the synthesized videos on three aspects: 1) Lip-sync Accuracy, 2) Image Quality, and 3) Video Realness. The average scores for each method are presented in Table 5. As shown, our method significantly outperforms the comparison methods in terms of Lip-sync Accuracy, indicating the effectiveness of the cross-modal alignment framework.
5.4 Model Inference Efficiency
We present the inference efficiency of different models in Table 6. The experimental results show that the proposed PASE achieves superior hardware efficiency while maintaining competitive inference speed. Specifically, PASE attains a 8.8× parameter reduction compared to Wav2vec 2.0 (37.7M vs 317M) with only 38.4% longer inference time (4.72s vs 3.41s), while significantly reducing GPU memory consumption by 57.3% (2051.5MB vs 4805.3MB). Compared to the lightweight Whisper, PASE shows comparable computational efficiency (4.72s vs 5.03s). These metrics collectively verify our design's effectiveness in balancing computational efficiency and phoneme-viseme alignment accuracy.
Model Inference Time (s) Params(M) GPU Usage (MB)
HuBERT [5] 3.36 $\pm$ 0.3 316 4348.2
DeepSpeech [4] 191.44 $\pm$ 3.7 - 523.2
Wav2vec 2.0 [6] 3.41 $\pm$ 0.3 317 4805.3
Whisper [7] 5.03 $\pm$ 0.2 19.8 1032.7
PASE(Ours) 4.72 $\pm$ 0.2 37.7 2051.5
Table 6: The results of the inference efficiency.
@@ 5.5 Ablation Study
To evaluate the effectiveness of the PASE framework, we conducted ablation experiments from two dimensions: speech spectrograms and modeling. The experimental results are presented in Table 7. As shown, the combination of STFT spectrogram and GRU significantly outperforms other variants in both video fidelity and lip sync accuracy, highlighting the necessity of the collaborative design of modules in PASE.
Speech Spectrogram Comparison. The STFT spectrogram exhibits a significant advantage over the Mel spectrogram. Under the NeRF rendering model, the LMD of STFT+GRU is reduced by 8.0% compared to Mel+GRU. This is because STFT’s linear frequency domain partitioning (0-8kHz full frequency range) preserves more frequency details. For example, STFT can retain high-frequency features of fricatives like /s/ and /S/ (4-8kHz), while the Mel (dominated by 0-4kHz) leads to the loss of such critical visual information. STFT spectrogram achieves the best or second-best performance across all variants, indicating that feature details are crucial for improving the performance of the rendering model.
@@ 6 Conclusion
To address the challenge of phoneme–viseme alignment ambiguity, we introduce PASE, a universal speech encoder that can be flexibly integrated into diverse talking portrait synthesis models. Unlike conventional approaches that align linguistic representations, PASE enforces phoneme-level cross-modal alignment between speech and lip shapes, guided by phoneme embeddings and contrastive learning. Furthermore, by combining STFT spectrograms with a GRU-based encoder, PASE effectively captures fine-grained temporal–frequency features. This design enhances robustness against noise and modality absence, while providing more discriminative and generalizable representations to support subsequent rendering models. Experimental results show that PASE achieves state-of-the-art performance on both NeRF and 3DGS rendering models. Notably, in terms of lip sync accuracy, PASE improves LSE-C and LSE-D by 13.7% and 14.2%, compared to the acoustic feature, and produces results close to the ground truth videos. Ablation experiments further present the effectiveness of the STFT-GRU combination.
7 Acknowledgements
This work is supported by the Natural Science Foundation of China (NSFC) under the grants No. 62572358, 621723064, and 62372334.